Electrolex, Inc., has four-year bonds outstanding that pay a coupon rate of 12.76 percent and make coupon payments semiannually. If these bonds are currently selling at $914.89. What is the yield to maturity that an investor can expect to earn on these bonds? (Round intermediate calculations to 4 decimal places, e.g. 1.2514 and final answer to 2 decimal places, e.g. 15.25%.) Yield to maturity % What is the effective annual yield? (Round intermediate calculations to 4 decimal places, e.g. 1.2514 and final answer to 2 decimal places, e.g. 15.25%.) Effective annual yield %
Answer #1

semi annual rate using excel function

=RATE(nper,pmt,pv,fv)

=RATE(4*2,1000*12.76%/2,-914.89,1000)

=7.8528%

yield to maturity =7.8528%*2=15.71%

the effective annual yield=(1+7.8528%)^2-1=16.32%